Calculate the amount in m years if a principal P is invested at a nominal annual interest rate of r compounded as given. Round to the nearest cent.
P=$480, m=4, r=7% compounded quarterly.
I have tried several times using info I got from this website to work this problem, but come up with some really far-fetched answers. Maybe I'm missing something along the way. Can anyone give me some specifics on how to set this up? Thanks!

Answer by stanbon(75887)

P=$480, m=4 yrs. , r=7% compounded quarterly.
A=P(1+r/4)^(4m)
A=480(1+0.07/4)^(4*4)
A=480(1.0175)^16
A=480*1.31992935
A=$633.57
Cheers,
Stan H.